OpenStudy (anonymous):

solve z^2+(1-i)z+(-6+2i)

OpenStudy (anonymous):

z^2+z-iz-6+26

OpenStudy (anonymous):

find the value of z in terms of i

OpenStudy (anonymous):

it is not an equation, there is nothing to 'solve'

OpenStudy (anonymous):

if it was for example \[z^2+(1-i)z+(-6+2i)=0\]then you could solve for \(z\) via the quadratic formula

OpenStudy (anonymous):

oh its =0 i forgot to write it! im trying to solve it by Q.F but not getting my ans

OpenStudy (anonymous):

by some miracle this one actually factors, as \[(z-2)(z+(3-i))=0\] but that is not so easy to see

OpenStudy (anonymous):

\[z=-\frac{(-1+i)+\pm\sqrt{(1+i)^2-4(1)(-6+2i)^2}}{2}\]

OpenStudy (anonymous):

all the work is finding \[(1-i)^2-4(-6+i)^2\]

OpenStudy (anonymous):

\[(1-i)^2=1-2i-1=-2i\]

OpenStudy (anonymous):

i dont get the underoot part

OpenStudy (anonymous):

you have to compute \[(1-i)^2-4(-6+2i)\] that is the hard work

OpenStudy (anonymous):

also why have you changed the signs?

OpenStudy (anonymous):

you get \[1-2i-1+24-8i\] \[=-10i

OpenStudy (anonymous):

24-10i

OpenStudy (anonymous):

changed the signs in front in the numerator because it is \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] and \(b=1-i\) so \(-b=-1+i\)

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

hmm i am thinking your job now is to take the square root of \(24-10i\)

OpenStudy (anonymous):

the answer is \(5-i\) but i am not sure how you are supposed to get it

OpenStudy (anonymous):

thats the main part on which im stuck since the last 2 hours

OpenStudy (anonymous):

i used wolfram the only other way i know how to do this is to write the number in polar form, then to it

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

or use a calculator that has complex numbers (most do now)

OpenStudy (anonymous):

mine doesnt